# Investigation of the Trailing-Edge Jet-Buffet Phenomenon Using Circulation Control at Transonic Speeds

### v1

@@ 4.2.2 Modal Decomposition
Figure  9 shows the normalized spatial amplitude $|\tilde{\Phi}|=|\Phi/\mathrm{max}(\Phi)|$ of both the first mode, which corresponds to the steady flow-field, and the first modal pair ($r=3,4$), which not only relates to the jet-buffet frequency but also has the highest modal amplitude among the buffet modes. As previously noted, increasing the jet momentum causes the upper surface shockwave to move further aft. From Fig. 9(a,c), it can be seen that while the upper surface shock is present in the first mode, it is absent in the modal pair and in subsequent modes. This absence is due to the jet’s inability to continue augmenting the circulation of the wing and as a result, cannot further shift the shock’s position. Therefore, it is understood that jet-buffet marks the end of circulation control, even prior to complete jet separation.
Also noted is a pressure structure seen on the surface of the trailing-edge at $r=3,4$ in Fig.  9d, which extends above the surface and contains a pressure core characterized by large gradients from its center, as well as a smaller pressure bubble further downstream. These structures match in form to the interaction of the trailing-edge shock with the unstable separation bubble.
@@
Superseding the latter modes are those shown in Fig. 10 in terms of their normalized spatial distribution. These modal pairs account for 44% of the modal influence and exhibit a maximal difference in modal amplitude of order $\mathcal{O}(10^{4})$. As the mode number increases, the spatial contribution decreases. The pressure structure first observed at $r=3,4$ becomes fragmented as the mode number increases, separating into an individual shock structure and a small pressure bubble. Meanwhile, its high-amplitude core steadily rises above the surface of the trailing-edge. The pressure bubble advances along the trailing-edge until it is shed into the wake, while the shock is gradually weakened and detached. When the higher modes are reached $r=11,12$ in Fig. 10e, the shock structure is completely detached, and the bubble is no longer visible. The initial formation and its fragmentation at increasing modal numbers suggest that the unsteady field is influenced by both the combined shock-pressure bubble and their individual components. The subsequent dissipation of both components indicates that above a certain mode number, the effects of the jet-buffet are diminished, and other effects, such as the shedding itself, become dominant. Furthermore, the initial structure observed in the first modal pair is no longer visible as we progress through the different modes. Therefore, it is understood that the fragmentation of the initial structure is the driving mechanism behind the jet-buffet cycle.
(a) global phase distribution.
(b) phase distribution at the trailing-edge.
Figure 11: Phase of the spatial distribution for $r=3,4$.
Additional information on the driving mode can be obtained by examining its phase shift $\phi_{\Phi}$, as shown in Fig. 11. Based on the phase contribution from the complex spatial distribution, i.e., $\Phi=|\Phi|e^{\phi_{\Phi}i}$, the analysis of the spatial phase allows us to pinpoint the direction of the pressure waves by tracing the shift from positive to negative values. Notably, in Fig. 11a, the phase shift indicates that the pressure wave moves from the leading-edge downstream to the trailing-edge, from the underside of the wing. Similarly, on the upper surface, the wrapping of the phase distribution around the leading-edge and the sudden bi-directional movement of the pressure waves some distance from it depict the upper surface shock. Regardless of its absence previously due its negligible spatial distribution magnitude. Further downstream, the presence of the trailing-edge shockwave compresses the pressure waves, resulting in a smaller spacing of the phase shift. When focusing on the trailing-edge (Fig. 11b), small regions are observed some distance away from its where the phase shift is reversed, indicating pressure propagation traveling upstream. Although the origin of this upstream propagation is difficult to ascertain from the phase distribution alone.
(a) $\tau=0$
(b) $\tau=0.27$
(c) $\tau=0.46$
(d) $\tau=0.54$
(e) $\tau=0.77$
(f) $\tau=1$
Figure 12: Time development of the pressure field, reconstructed using the first modal pair $r=3,4$.
@@
Based on the previous observation, the time development and contribution to the pressure reconstruction of the driving mode, i.e., $p_{3}=b_{3}\Phi_{3}e^{\lambda_{3}t}$, can be examined, as shown in Fig. 12, for different pressure distributions along the jet-buffet cycle. Where $\tau=0,1$ correspond to the reconstructed maxima of the single mode. During these times, the high amplitude of the shock-bubble structure on the trailing-edge has a positive sign due to the jet attachment. Meanwhile, at $\tau=0.46,0.54$, the combined structure has a negative sign due to the jet detachment. Additionally, $\tau=0.27,0.77$ represent the points in the cycle where the trailing-edge pressure structure shifts its phase. As further illustrated in Fig. 13 it is observed that when the phase shifts, the pressure bubble is in anti-phase with the shock formation which subsequently separates from the trailing-edge surface. Thus, the driving mechanism of the jet-buffet relies on the feedback from the bubble downstream. As the cycle progresses, pressure waves propagate from the pressure bubble to the shock formation and vice versa, causing an interference that detaches and reattaches the shock to the surface, thus creating the jet-buffet.
(a) $\tau=0.27$
(b) $\tau=0.77$
Figure 13: Reconstructed pressure using $r=3,4$ at the temporal location where phase shift occurs between the pressure bubble and the shock formation.
@@ 5 Conclusions
A numerical validation via RANS was performed for an experimental study investigating the applications of circulation control (CC) on an elliptic airfoil at transonic conditions. The steady-state results aligned well with the experimental study, matching the pressure distribution and capturing the major effects, such as the upper surface shock and trailing-edge separation bubble, albeit with slight inconsistencies in their locations.
Upon assessing the contribution of CC at higher momentum coefficients ($C_{\mu}=0.01$), unsteadiness was observed, as seen in previous studies, prompting the use of a URANS simulation to investigate its impact on the flow-field. The analysis revealed an interaction between the high-momentum jet, the trailing-edge shock, and the separation bubble, which caused the bubble to disperse. This interaction induced oscillations in the lift coefficient, a phenomenon referred to as the jet-buffet. Further investigation through Dynamic Mode Decomposition (DMD) revealed that the jet-buffet dynamics were largely decoupled from the upper surface shock. This decoupling resulted from the inability of the jet to further augment the wing’s circulation, preventing the wing from conforming to super-circulation.
The origin of the jet-buffet was explored through the dynamic modes obtained via DMD analysis. It was determined that the phenomenon is driven by the formation of a shock-bubble structure, with a distinct pressure core at its center. As the modes progressed, the structure fragmented, producing higher-order modes in which the pressure bubble moved along the trailing-edge before being shed into the wake.
The time development of the driving mode further elucidated its origins. As the cycle progressed, pressure propagation between the shock and separation bubble caused interference, resulting in shock detachment from the surface. This mechanism was responsible for both the detachment and subsequent reattachment of the jet.
The implications of these findings suggest that while the jet-buffet leads to the degradation of the CC mechanism under these conditions, it could also result in structural oscillations and impose a significant cost to the wing’s stability due to rapidly changing wing moments. Thus, the presence of the jet-buffet could influence the practical application of CC for transonic flight.

### v3

@@ 4.2.2 Modal Spatial Distribution
Figure  14 shows the normalized spatial amplitude $|\tilde{\Phi}|=|\Phi/\mathrm{max}(\Phi)|$ of both the first mode, which corresponds to the mean pressure field, and the first modal pair ($r=3,4$), which not only relates to the bi-stability frequency but also has the highest modal amplitude among the buffet modes. The second mode $r=2$ as discussed previously, possessed almost no modal influence, and did not contain additional information from its spatial distribution which was found to be relevant to the bi-stability. As previously noted, increasing the jet momentum while the bi-stability persists, does not change the location of the upper surface shockwave. From Fig. 14(a,c), it can be seen that while the upper surface shock is present in the first mode, it is absent in the modal pair and in subsequent modes. This absence, which was outlined in the URANS results, is due to the jet’s inability to continue augmenting the circulation of the wing and as a result, cannot further shift the shock’s position. Therefore, it is understood that bi-stability marks the end of circulation control, even prior to complete jet separation. Also noted is a pressure structure seen on the surface of the trailing-edge at $r=3,4$ in Fig. 14d, which extends above the surface and contains a pressure core characterized by large gradients from its center, as well as a smaller pressure bubble further downstream. These structures match in form to the interaction of the trailing-edge shock with the unstable separation bubble.
@@
Superseding the latter modes are those shown in Fig. 15. These modal pairs account for 44% of the modal influence, and as the mode number increases their spatial contribution decreases. The pressure structure first observed at $r=3,4$ becomes fragmented as the mode number increases, separating into an individual shock structure and a small pressure bubble. Meanwhile, its high-amplitude core steadily rises above the surface of the trailing-edge. The pressure bubble advances along the trailing-edge until it is shed into the wake, while the shock is gradually weakened and detached. When the higher modes are reached $r=11,12$ in Fig. 15e, the shock structure is completely detached, and the bubble is no longer visible.
From the perspective of the system’s damping characteristics, modes r=36r=3-6 as well as $r=13,14$ are underdamped ($0<\zeta<1$) and thus contribute to the stable oscillatory portion of the unsteady flow. In contrast, modes r=712r=7-12 are classified as driving modes ($\zeta<0$) and are responsible for destabilizing the system. The difference in spatial structure between these two groups offers insight into this contrast: modes r=36r=3-6 preserve the initial pressure configuration, exhibiting only slight displacements of the pressure core as the structure begins to deform. Conversely, modes r=712r=7-12 reveal a fragmented pressure field, indicating their role in the breakup process. Consequently, it is this breakup of the pressure structure that introduces destabilizing dynamics into the system.
(a) $r=15,16$
(b) $r=19,20$
(c) $r=23,24$
(d) $r=27,28$
(e) $r=29,30$
(f) $r=33,34$
Figure 16: Spatial distribution of positive damping modes related to trailing-edge vortex shedding.
Modes $r=13,14$ are also an underdamped modal pair; however, unlike the earlier underdamped modes, they exhibit substantial dissipation of both the shock and the pressure bubble. This extensive dissipation suggests that beyond a certain mode number, the influence of the bi-stability phenomenon weakens, while other dynamics, such as vortex shedding, begin to dominate. This transition is illustrated in Fig. 16, which presents the spatial structure of the underdamped modes following $r=13,14$. Although these higher-order modes still capture the detached shock seen in $r=13,14$, they differ in the near-wake region, where vortex shedding becomes prominent across all modes. Furthermore, the two-order-of-magnitude difference in amplitude between the lower underdamped modes associated with the initial bi-stability structure (r=36r=3-6) and the higher-order shedding modes highlights a gradual shift from the bi-stability mechanism to vortex shedding. As a result, the latter modes contribute minimally to the bi-stability. The absence of the coherent structures observed in the first and second modal pairs reinforces the conclusion that those early modes are the primary governing mechanism of the bi-stability cycle.
4.2.3 Mode Time Development
Figure 17: Time history of $p_{3,4}$ (corresponding to $r=3,4$)
depicting the interaction and feedback of the trailing-edge shock wave with the upstream pressure bubble.
@@ 4.2.3 Mode Time Development
To delineate on the observation, the time development and contribution to the pressure reconstruction of the underdamped modal pairs, $p_{3,4}$ and $p_{5,6}$ where, $p_{j}=b_{j}\Phi_{j}e^{\lambda_{j}t}$, can be examined. Shown in Fig. 17 is the time development of $p_{3,4}$, for different pressure distributions along the bi-stability cycle. Where $\tau=0,1$ correspond to the reconstructed maxima of the single mode. During these times, the high amplitude of the shock-bubble structure on the trailing-edge has a positive sign due to the jet attachment. Meanwhile, at $\tau=0.45,0.52$, the combined structure has a negative sign due to the jet detachment. Additionally, $\tau=0.27,0.75$ represent the points in the cycle where the trailing-edge pressure structure shifts its phase. At these temporal locations, it is observed that when the phase shifts, the pressure bubble is in anti-phase with the shock formation which subsequently separates from the trailing-edge surface. Thus, the primary mechanism of the bi-stability relies on the feedback from the bubble downstream. As the cycle progresses, pressure waves propagate from the pressure bubble to the shock formation and vice versa, causing an interference that detaches and reattaches the shock to the surface. This finding aligns with the experimental conclusions for the detachment of a supersonic jet at quiescent air [25] where the bubble played a major role in the bi-stable mechanism of attachment and reattachment. Nevertheless, unlike the bi-stability at quiescent air, the absence of a dominant anti-phase with another pressure bubble, solidifies the difference between the the two cases. The presence of the trailing-edge shock plays a major role in the stable periodic motion of the bi-stability.
Figure 18: Time history of $p_{5,6}$ (corresponding to $r=5,6$)
depicting the two distinct phase shifts of both the trailing-edge shock and pressure bubble.
Since all modes are linearly correlated, the second modal pair ($r=5,6$) corresponds to the second harmonic of the primary mode. As illustrated in its pressure reconstruction in Fig. 18, two distinct phase shifts can be identified: one at $\tau=(0.16,\,0.40,\,0.65,\,0.90)$ and another at $\tau=(0.03,\,0.28,\,0.53,\,0.78)$. The first set reflects the separation of the combined pressure structure observed in $r=3,4$ into two discrete structures, enclosing the pressure bubble and interacting through a feedback mechanism. The second set is more subtle, matching to a similar reconstruction process but focused on the pressure bubble itself, with the shockwave remaining static. Given that the modal pair $r=5,6$ was previously associated with the reshaping of the initial pressure structure, these phase shifts must correspond to feedback processes responsible for that reshaping. Thus, the second modal pair is not responsible for initiating attachment or detachment, but instead characterizes how the pressure field reorganizes as a result of those processes.
4.2.4 Pressure Field Reconstruction
The conclusions from prior sections, are used as a foundation for the pressure reconstruction. Since the $r=3,4$ and $r=5,6$ represent the underline detachment, reattachment and pressure redistribution mechanism, if added to the averaged pressure field, should capture the bi-stability mechanism and serve as a reduced order model for the effect. To determine this, the error in prediction is evaluated by the root mean square of the pressure field, $\mathrm{RMSE_{p}}$ at each cycle time, normalized by the dynamic pressure, to create a representation of the deviation in the pressure coefficient, i.e. $\delta C_{p}=\mathrm{RMSE_{p}}/Q_{\mathrm{dyn}}$.
Figure 19: Accuracy of the pressure field reconstruction for both the total number of modes as well as up to the six most contributing modal pairs.
$r=1$, r=13r=1-3, r=14r=1-4, r=15r=1-5, r=16r=1-6
Figure 19 shows the reconstruction error as a function of increasing modal number. The error decreases rapidly as more modes are included. Using only the first mode results in a 10% error, but the reconstructed signal remains entirely static. In contrast, reconstruction using the first 24 modes reduces the error to 0.5%, albeit at the cost of including a larger number of terms in the reduced model. This trend aligns with the earlier observation from the modal influence analysis, where most of the flow’s dynamic content was carried by a small subset of modes.
Conversely, when additional modes are included ($r>24$), a persistent error peak emerges near the end of the cycle. The consistent location of this peak, and its appearance only beyond a certain mode count, suggests that a specific dynamic behavior arises at this stage of the cycle that is not adequately captured by the DMD. It may stem from the transition of the flow from separated shear-layer and vortex shedding back to bi-stability. Nevertheless, the maximum 2% error associated with it remains within typical engineering tolerances.
Figure 19 also illustrates this behavior in more detail, showing the temporal evolution of reconstruction error for the lower modes. With only the first mode, the error profile exhibits an harmonic trend, indicating a failure to capture the underlying oscillatory behavior. As more modes are added, the error profile flattens, with only small residual oscillations. When all six significant modes are included, the maximum error drops to approximately 2%, occurring only near the cycle edges. These results underscore not only the importance of the $r=3,4$ and $r=5,6$ modal pairs, but also the ability of a compact ROM to represent the complex unsteady pressure field and its associated lift dynamics with up to 98% accuracy.
@@ 5 Conclusions
A numerical validation via RANS was performed for an experimental study investigating the applications of circulation control (CC) on an elliptic airfoil at transonic conditions. The steady-state results aligned well with the experimental study, matching the pressure distribution and integral coefficients.
Upon assessing the contribution of CC at higher momentum coefficients ($C_{\mu}=0.01$), unsteadiness was observed, as seen in previous studies, prompting the use of a URANS simulation to investigate the unsteady flow-field. Both a quantitative comparison with the TDT time-averaged data, as well as a visual comparison with the similar effects, revealed an interaction between the high-momentum jet, the trailing-edge shock, and the separation bubble.
This process caused the bubble to be pushed beyond the trailing-edge tip and thus destabilize the jet. This interaction induced bi-stable oscillations in the lift coefficient. Further investigation revealed that the bi-stability dynamics were largely decoupled from the upper surface shock. This decoupling resulted from the inability of the jet to further augment the wing’s circulation, preventing the wing from conforming to super-circulation.
To both understand the feedback mechanism necessary to create persistent oscillations, as well as deduce if the bi-stability can be approximated by a reduced model, the bi-stability was explored through the use of Dynamic Mode Decomposition (DMD). It was determined that the phenomenon is driven by primarily by the first two modal pairs which encapsulate 24.5% of the modal influence are associated either with the primary pressure structure, or its initial redistribution. With higher modes leading to the its fragmentation or in its transition vortex shedding.
The time development of these modes further elucidated on the effect. For the first modal pair, as the cycle progressed, pressure propagation between the shock and separation bubble caused interference, resulting in shock detachment from the surface. This mechanism was responsible for both the detachment and subsequent reattachment of the jet. The second modal pair complemented the first by an additional feedback mechanism that drives the pressure redistribution.
These findings reveal that under the examined conditions, bi-stability fundamentally disrupts the circulation control mechanism. Yet, despite its complexity, the phenomenon is accurately captured using URANS and distilled even further into a ROM model. Remarkably, only six dynamic modes are sufficient to reconstruct the unsteady pressure field with 98%, offering a powerful, efficient path toward its modeling